Calculate the force of attraction between two small objects of mass 5kg and 8kg respectively which are 10cm apart. (G = 6.7 x 10^(-11) Nm^2kg^-2. Comment on the value of your answer.
1
Expert's answer
2021-02-10T10:10:52-0500
"F=G\frac{mM}{r^2}\\F=(6.7\cdot10^{-11})\frac{5\cdot8}{0.1^2}=2.7\cdot10^{-7}N"

Very small force.